OpenStudy (anonymous):

This will be my last question for this evening. Write solve the equation: 51/2%(five an a half percent) of what number of cases is 33 cases.

OpenStudy (anonymous):

600

OpenStudy (anonymous):

is this d ans

OpenStudy (anonymous):

Tks sweetznchillz but i keep coming out with 6000?

OpenStudy (anonymous):

{(11/2)/100}*x=33 (11/200)*x=33 x=(33*200)/11 x=3*200 =600

OpenStudy (anonymous):

sweetznchillz, i see where i went wrong. Thank you